Graph the surface area of a cube as a function of the volume of a cube


1
Expert's answer
2022-04-07T13:23:45-0400

If the side of a cube is x then the surface area of a cube is "S=6x^2" and the volume is "V=x^3".

Then "x=V^\frac13" and "S=6x^2=6V^\frac23" .

The surface area of a cube as a function of the volume of a cube:

"S(V)=6V^\frac23"